we are conducting many hypothesis tests to test a claim. Approximately how many of the tests will yield significant results by chance, if the null hypothesis is true?

a. 600 tests using a significance level of 5%.

b. 70 tests using a significance level of 10%.
Answer #1

a) Number of tests that will yield significant results by chance = 600*0.05 = 30

b) Number of tests that will yield significant results by chance = 70*0.10 = 7